OpenStudy (anonymous):

suppose A is a (3x1) nxm matrix such that A= (-4) (0) (0) = (3x1) matrix: (36) (12) (-4) what is the first column of A?

OpenStudy (anonymous):

your question isnt really clear, perhaps you could use the LaTex equation writer in a comment to write the matrices out properly? then i can help.

OpenStudy (anonymous):

suppose that A is an nxm matrix such that A= \[\left[\begin{matrix}-4 \\ 0\\0\end{matrix}\right]\] = \[\left[\begin{matrix}36 \\ 12 \\-4\end{matrix}\right]\] what is the first column of A?

OpenStudy (anonymous):

im sorry, i cant see how that works, how can A be both of those things?

OpenStudy (anonymous):

here's a screenshot of the problem to make it more clear

OpenStudy (anonymous):

aaah i understand now, A is multiplied by the first matrix to get the second

OpenStudy (anonymous):

let\[A= \left( \begin{array}{ccc} a & b & c \\ d & e & f \\ g & h & i \end{array} \right)\] now do you know how matrix multiplication works?

OpenStudy (anonymous):

matrices are only compatible for multiplication if this is true: if matrix A is n x m and matrix B is p x q , then m = p also the product of A and B will be of order n x q therefore we can deduce A must be 3 x 3

OpenStudy (anonymous):

ok, thank you